OpenStudy (lilkg77):

The first quartile of a data set is 72, and the third quartile is 92. Which of these values in the data set is an outlier? A. 121 B. 101 C. 41 D. 61

Directrix (directrix):

Subtract the lower quartile 72 from the upper quartile 92. Post what you get. The number you get is known as the Interquartile Range, denoted by IQR.

Directrix (directrix):

@lilkg77 ^^^ Please answer so that we can go to the next step. Thanks.

OpenStudy (lilkg77):

20

Directrix (directrix):

Okay. What is 1.5 times 20 =

Directrix (directrix):

@lilkg77 ^^^ Question waiting for you. :)

OpenStudy (lilkg77):

30 @Directrix

OpenStudy (lilkg77):

@Directrix

Directrix (directrix):

Correct. A data point is an upper outlier if it if has value greater than the upper quartile + 30. Q3, the upper quartile, is 92. Q3 + 1.5 IQR = 92 + 30 = 122. Look at the answer options. Is there a number given which is greater than 122. @lilkg77

Directrix (directrix):

After you answer, we will check to see if there is a lower outlier.

Directrix (directrix):

>>Is there a number given which is greater than 122. Choose from these: A. 121 B. 101 C. 41 D. 61

Directrix (directrix):

@lilkg77